Find two consecutive integers such that half the smaller plus three times the larger is -32
Answer by stanbon(57203)

Find two consecutive integers such that half the smaller plus three times the larger is -32
---------------------
1st: x
2nd: x+1
---------
Equations:
(1/2)x + 3(x+1) = -32
--------
Multiply thru by 2 to get:
x + 6(x+1) = -64
7x + 6 = -64
7x = -70
x = -10
============================
1st: -10
2nd: -9
===============
Cheers,
Stan H.